OpenStudy (anonymous):

A skier of mass 40 kg, pushes off the top of a hill with an initial speed 4.0 m/s. Neglecting friction, how fast will she be moving after dropping 10 m in elevation?

OpenStudy (anonymous):

is there any angle of inclination given???

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i figure no friction means acceleration due to gravtiy which is almost 10m/s^2. So the skier has probably only been moving for about a second but idk because of the 4m/s

OpenStudy (anonymous):

so take it to be 0 degrees from the horizontal

OpenStudy (anonymous):

so vertical initial velocity is zero so apply v=u + at

OpenStudy (anonymous):

what does the u represent?

OpenStudy (anonymous):

initial velocity = 0

OpenStudy (anonymous):

also there is no time. are we just assuming t= 1 because she has moved 10 m

OpenStudy (anonymous):

oh sorry use v^2 = u^2 + 2ah

OpenStudy (anonymous):

so it would be -14m/s correct?

OpenStudy (anonymous):

or 14 m/s

OpenStudy (anonymous):

yup.... magnitude is 14 the sign just tells u d direction which is ur convention

OpenStudy (anonymous):

yes thank you

OpenStudy (anonymous):

u r welcum